1) A confidence interval for a population proportion p is found to be (0.32, 0.36). The margin of error is:

2) A confidence interval for a population proportion p is found to be (0.32, 0.36). The sample proportion ?̂ is:

3) Given Ha: p ≠ 0.85 and α = 0.05, which level of confidence should you use to construct a confidence
interval that corresponds to this hypothesis test?
Answer #1

given, CI is (0.32, 0.36)

sample proportion, = (0.32+0.36)/2 = 0.34

Lower Confidence Level = -Margin of error

M.E = 0.34-0.32 = 0.02

3) =0.05

1- = 0.95

i.e., 95% level of confidence should be used to construct the confidence interval